OpenStudy (anonymous):

By solving the equation x^3+1=0, find the three cube roots of -1. If one of the complex cube roots is A express in terms of A. Prove that 1+A^2=A.

OpenStudy (anonymous):

Without using complex roots?

OpenStudy (anonymous):

using complex roots

OpenStudy (anonymous):

Using this \[\Large a+bi=r(\cos\theta \ + \ i\sin\theta)\]form?

OpenStudy (anonymous):

why bother

OpenStudy (anonymous):

Yeah, why bother :D We can factor x^3 + 1 as it's the sum of two cubes :)

OpenStudy (anonymous):

\[x^3+1=(x+1)(x^2-x+1)\] quadratic formula gives the second two zeros

OpenStudy (anonymous):

could you be more specific?

OpenStudy (anonymous):

\[\Large x^3+1=(x+1)(x^2-x+1)\] Which means -1 is one of the roots of the equation right?

OpenStudy (anonymous):

So the other two roots are whatever the roots of this equation is... \[\Large x^2-x+1=0\] Easily obtainable through the quadratic formula :)

OpenStudy (anonymous):

ok, thx